OpenStudy (anonymous):

What is the solution of the system? 2x+6y=14 and -x+2y=3

OpenStudy (anonymous):

\[2x+6y = 14\] \[-x + 2y = 3\] and you can rewrite the first one: 2x + 6y = 14 but you can multiply 2 to the both sides of the second one: 2(-x+2y) = 2x3

OpenStudy (anonymous):

got it?